Find the simplest form for the following boolean expressions using algebraic methods.
1. (A+B̄)(B+C)+(A+B)(C+Å̄)
2. (A+B)(AC+AC̄)+AB+B
1
Expert's answer
2020-06-09T18:36:39-0400

1 )


"AB+AC+B.\tilde{B}+\tilde{B}.C+AC+A.\tilde{A}+BC+B.\tilde{A}"

"AB+AC+\tilde{B}.C+AC+BC+B.\tilde{A}"

"AB+AC+\tilde{B}.C+BC+B.\tilde{A}"

"AB+AC+C(\tilde{B}+B)+B.\tilde{A}"

"B(A+\tilde{A})+AC+C"

"B+AC+C"

"B+C(A+1)"

"B+C"







2 )

"(A+B)(A.C+A.\tilde{C})+A.B+B"

SOLUTION:-

="A.C+A.\tilde{C}+A.B.C+A.B.\tilde{C}+A.B+B"

="A.C+A.\tilde{C}+A.B.C+A.B.\tilde{C}+A.(B+1)"

="A.C+A.\tilde{C}+A.B.C+A.B.\tilde{C}+A"

="A.C+A.\tilde{C}+A.B.(C+\tilde{C})+A"

="A.C+A.\tilde{C}+A.B+A"

="A.C+A.\tilde{C}+A.(B+1)"

="A.C+A.\tilde{C} +A"

="A.(C+\tilde{C})+A"

=A+A

=A


LAWS:-

1 ) "x.\tilde{X}=0"

2 ) x.x=x

3 ) x+x=x

4 ) x+1=1

5 ) "x+\tilde{x}=1"